Home‎ > ‎Reports‎ > ‎Crew098‎ > ‎

day11


Crew 98 Log Book for January 25, 2011

Photos of the day







Commander's Check-In

Virgiliu Pop Reporting

Crew Physical Status: Excellent


Time Departed/Returned from EVA: 4PM-6PM


Brief Narrative of Field Mission Results: Rover (GRIVEI) mission.


EVA Data/Interpretations: As per the EVA report.


Engineering/Hab Maintenance: Another visit from DG, the friendliest Martian ever.


Report Transmission Schedule: Intended by 8 PM.


Plans for Tomorrow: Will be decided in the morning.


Inventory: One Motorola radio missing; we will order a replacement from Amazon.com


Support Requested: -


Miscellaneous: -


Commander's Report

Virgiliu Pop Reporting

(NOTE FROM MISSION SUPPORT: Sept 19th is International Talk Like a Pirate day... but obviously that holiday was today on Mars. This could be the start of something big!)


Avast, ye land-lubbers! Here be today's report! Arrr, the Captain is be'arrr'ded, he looks now like a Pirate, and while today it was not a cookin' day, he be known as Captain Blackbeard.


Fun may be, yet he and the members o' the the crew 'arrr' highly disciplined in their exploration o' M'arrr.'s He likes his crew t' work h'arrr'd and play h'arrr'd; and work they did, shi'er me timbers.


Aye, H, with help from the friendliest and most helpful local resident, DG, tended t' her sextant, while the other two lady bucaneers analised the pieces o'eight (read samples) from yesterday's E'as. The captain then took the other two bucaneers - H and D - in a most ad'enturous Eva, o'er se'en seas, and D then took many a picture. Gar, what a booty!


Arrr, at sunset the away team returned t' the ship with his crewmen and the three din'hies, and now is waitin' for I and B' cook tonight's meal, which will surely protect them from scur'y. Ahoy, me hearties!


Engineering Report

Haritina Mogosanu Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): used

Honey (Gas Generator) (Used? Yes/No): not used

Battery Charge Level:?
29.5 V everstart
34.4 V armor plate
40.3 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
1 generator run all day ok


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty):1/2

Propane Tank (Full/0.75/0.5/0.25/Empty):30%

Gasoline Tank (Full/0.75/0.5/0.25/Empty):1 containers full x 5 gallons

Notes/Comments:


Oil: not checked DG said to leave it on him

Kitty Oil Quantity (# of Quarts in storage):2.5 gal

Honey Oil Quantity (# of Quarts in storage):3 gal

ATV Oil Quantity (# of Quarts in storage):on the engineering work bench
2 x 946 ml

Car Oil Quantity (# of Quarts in storage):where is it?


Water Status:

Outside Potable Water Tank Level (inches from bottom): 11

Trailer Potable Water Tank Level (inches from bottom): almost full

Hab Potable Water Tank Level (inches from bottom): 12

Potable Water Meter Reading: 412841

Notes/Comments:


GreenHab:

Condition of plants in Tank 1: plants dead

Condition of plants in Tank 2: very poor

Condition of plants in Tank 3: improving

Greenhab notes/ comments:


Observatory:

Telescope Used? (Yes/no) yes

Observatory Notes:

Telescope underwent realigning and reinforcing as follows:

We removed the telescope tube off the dove tail and placed it securely on the ground.
DG cut the faulty bolt and then removed the mount.
The inside part of the BOLT was broken and DG welded a new one.
We measured north by Polaris (as instructed by Peter) and double checked it with the magnetic compass, computed the variation of 14E (data taken from Hanksville airport) of it.
It looked that if we only pulled the main pier sideways a little bit we would be able to get the orientation so we decided to give it a go without drilling new holes. (this was also based on the fact that after looking inside the pier upon opening it, DG realised that he did not have the proper tools for redrilling properly aligned bolts. )

We placed the electonics pier and the counterweights and strenghtened the bolts back on the main pier.
We put the telescope back on the dove -tail, balanced it in RA and Dec.
We leveled the pier.

Tonight I'll polar align it and see how it tracks etc.
Please keep all your fingers crossed!


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no):Yes
Oil Checked (yes/no):No
Tire Status:not checked
Odometer at end of day:-123233
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: y/n y
Oil Checked: y/n y
Fuel Consumed: filled
Tire Status:checked
Comments:

ATV 2 (Spirit):
Used: y/n y
Oil Checked: y/n y
Fuel Consumed:replaced
Tire Status:checked
Comments:

ATV 3 (Viking I):
Used: y/n n
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status: not checked
Comments: started

ATV 4 (Viking II):
Used: y/n y
Oil Checked: y/n y
Fuel Consumed: replaced
Tire Status: checked
Comments:


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 75

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 65


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: cleaned the glass on the front of the camera and reset the computer.


General Engineering
1. Notes/Comments: DG was here again and made major repairs to the plumbing and the insulation as you may know.

2. We bought a new green laser pointer that we wish to donate for the Musk Observatory. This will come in handy for both visual observing as well as alignments of the telescope.

3. The hand control of the telescope has been previously dropped on the floor. I am basing this affirmation on the fact that when i first used it, the scrren was croocked and I was mentioning in a previous email to Peter that the display went off; DG brougt a new adaptor as we thought that happened due to power fluctuations. Last night the screen went off again and came back to life after I pressed on it so I tightened up the screws. It works for now but it needs looking at it and glueing back the LCD screen.

EVA 17 Report

Haritina Mogosanu Reporting

Time: 16.00-18.30


Crew Members: Haritina Mogosanu, Dragos Bratasanu, Virgiliu Pop


Site Location: 38.454039,-110.790139


Transit Mode: ATV


Objectives: test drive the rover; photo opportunity ;


Results: Rover succesfuly tested on rough terrain, photos taken


Lessons Learned: it was very hard to breathe and walk on the sloppy terrain, the visibility was not too good either as it's hard to be astute when you don't see where to step. We decided that is also very hard to drive with the back packs on the ATVs so we secured them with ropes at the back.


Cook's Report

Beatrice Gilea Reporting

1a) Today was
- a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?
- No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.
- Texas BBQ Chicken with Beans

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)
- 2 servings Texas BBQ Chicken with Beans discarded

4) Recipes: - None

5) Comments/questions for the food study investigators
- The crew didn't like the food, except Virgil and Dragos who found it acceptable